OpenStudy (nerdgirl):

HELP!!!! Will medal and fan!! (Yes, I know this is the wrong subject but no one's answering me) Tin reacts with fluorine to form two different compounds, A and B. Compound A contains 38.5 g of tin for each 12.3 g of fluorine. Compound B contains 56.5 g of tin for each 36.2 g of fluorine. What is the lowest whole-number mass ratio of tin that combines with a given mass of fluorine?

OpenStudy (anonymous):

A contains 38.5 g of tin for each 12.3 g of fluorine: mole ratio: (38.5 g)/(118.71 g/mol):(12.3 g)/(18.998 g/mol) = 0.324:0.647 = 1:2 ⇒ SnF₂ B contains 56.5 g of tin for each 36.2 g of fluorine: mole ratio: (56.5 g)/(118.71 g/mol):(36.2 g)/(18.998 g/mol) = 0.476:1.905 = 1:4 ⇒ SnF₄

OpenStudy (nerdgirl):

So what's the final ratio?

OpenStudy (nerdgirl):

My Chemistry teacher said 2:1... how did he get that?

OpenStudy (anonymous):

21

OpenStudy (nerdgirl):

alright how did you get it? That's what I don't understand.

OpenStudy (anonymous):

you add 9+10

OpenStudy (nerdgirl):

@Riamarializz can you tell me how my teacher got 2:1?

OpenStudy (nerdgirl):

Here: compound A: 38.5 Sn/12.3 g F = 3.12 g Sn/1.00 g F Compound B: 56.5 g Sn/36.2 g F= 1.56 g Sn/1.00 g F, so 3.12 g Sn (compound A) / 1.56 g Sn (compound B) = 2:1

OpenStudy (nerdgirl):

This is the answer my teacher gave me. Can someone help me make sense of it?

OpenStudy (nerdgirl):

@thomas5267

OpenStudy (thomas5267):

Busy for now. Will answer later.

OpenStudy (nerdgirl):

great...